OpenStudy (anonymous):

Darnell is selecting names from a hat to determine who will choose the family activities for the weekend. The names in the hat include his mother, father, sister, and him. Calculate the probability of each compound event. Match Term Definition (Darnell and sister), without replacement (mother and father), with replacement (parent and sister), with replacement (parent and child), without replacement a1/3 b1/16 c1/12 d1/8

OpenStudy (anonymous):

\[P(D, \S) = 1/4*1/3=1/12\] \[P(M,F) = 1/4*1/4= 1/16\]

OpenStudy (anonymous):

\[P(P,\S) =2/4*1/4 =1/8\] \[P(P,C)=2/4*2/4=1/4\]

OpenStudy (anonymous):

thanks